To: J. Albertine, D. Helms, W. Skinner
cc: Division (M7), FEL Coordination Group
From: F. Dylla
Subject: IRFEL Weekly Report, September 8-12, 1997
Date: September 12, 1997
Installation and commissioning preparations dominate the week's
activities. Highlights include installation of the back leg vacuum
system, the majority of the arc magnet stands, and three of the
optical chicane dipoles. In addition, bakeout of the gun vacuum
system began in preparation for the gun re-commissioning.
On Thursday, Sept. 11, Bob Marianelli and Eric Rohlfing from the
DOE-BES Program Office, and Dan Lehman from the DOE-ER Construction
Management Office visited the laboratory to review the progress
on the IR Demo project and discuss future DOE interactions. Eric
Rohlfing is reviewing the BES pre-proposals in Chemical Sciences
that were submitted to DOE last week. BES management requested
the project commissioning schedule and a listing of current and
potential university collaborators (particularly at the local
universities). We also discussed the options available for extending
the IR Demo wavelength range. Of particular interest to BES is
the mid-to-far IR range because of the scientific applications
noted in the 1994 NAS Study. Note: that the IR Demo with the
present 42 MeV linac can deliver long wavelength radiation to
approximately 25 mm. A brief white
paper on the wavelength upgrade options available to the IR Demo
will be prepared for BES.
H. Grunder and F. Dylla made a presentation on Sept. 10 to SURA
Executive Committee on the status of FEL program. A related discussion
concerned possible involvement of the SURA Materials Council with
FEL User Lab proposals and sponsorship of a proposed workkshop
on Laser Processing in January 1999.
Preparations were made for next week's SURA/DOE Science and Technology
Review. Review team members who will be paying particular attention
to the FEL Program presentation are John Galayda (Advanced Photon
Source), Erik Johnson (BNL) and Steve Laderman (Hewlett Packard
Financial data was collected this week for preparation of the
project monthly report for August.
Regarding the beam-transport system:
After a series of at out-of-plane measurements of the integral
field across the 9 cm good field region that validated D. Douglas'
assumptions, all eight optical chicane dipoles (DWs) were signed
off. Three were installed. The standard DW was held back to
perform a representative field map. In addition, measurements
were taken using 1006 steel field clamps substituting for the
production 1018 clamps. Virtually identical field integrals were
The first DU magnet was measured using the hall probe with continued
measurements on it and the first article DV planned for early
next week. The remaining seven magnets of this series were shipped
The final configuration of the field clamp on the DV dipole at
the intersection with the 10 MeV lines (and the configuration
of any magnetic shielding) had been left to the magnet measurement
phase of the magnet's commissioning. However, our tests indicate
this approach is not prudent. Field clamps have proved to be
exceptionally influential in maintaining uniform field integrals
across the good field regions and cut-and-try modification will
not converge on a solution fast enough. We will rather define
the configuration using finite element magnet modeling using TOSCA.
Tom Schultheiss from Northrop Grumman is consulting next week
with Robin Wines, bringing a TOSCA model of the DW dipole that
he can morph into the DV style to speed up the analysis.
Assembly of the eight reverse bend dipoles at Everson Electric
continued with shipping expected around September 24.
Arrangements with Everson to assemble the 180 Degree Dipoles are
nearly complete and the cores and parts were shipped to them from
Process Equipment Co. However, gluing the segmented mu metal
and brass to these cores will use bolts to provide the flattening
force rather than the press used on previous magnets. Everson
wants several trials on small prototypes first to prove this
different process. The resulting delay may stretch final assembly
of a first magnet to end of October. We are proposing to do a
trial assembly of the second magnet's parts first with a non glued
substitute for the mu metal and brass. The assembly also would
prove out the coil placement and all magnet hardware. We would
ship it here and install that assembly in late September into
the first arc to prove out the stands, vacuum chamber for the
complete arc and the diagnostics that mount at the 90 degree point
in the magnet. Final assembly of the first magnet would proceed
In the quadrupole and sextupole area, winding of coils and manufacture
of the cores for the eight Panofsky trim quads continued (with
about 60% of the task done) at Magnet Enterprises International
In the corrector area, manufacture of the eight coils for the
phasing dipoles (DGs) continued at MEI. Mu metal cores as well
as the cores for the other reverse bend correctors (DF&DCs)
are being fabricated at Magnetic Shield Corp. and fabrication
of the 8 coils for the DFs and 8 coils for the DCs was started
at a new vendor, Advance Magnet Lab.
Bids for air core corrector half-coil sub-assemblies are due in
September 15. In house manufacture of one additional prototype
horizontal air core corrector (DB) and two inner vertical correctors
(DH) continued. Magnetic measurements of the rough prototype
indicated total harmonic content on a 2 inch radius of 2% with
a 1% quadrupole dominating. Tests will continue to see if the
quadrupole can be nulled out by adjusting the relative placement
of the half coils.
All reverse bend chamber bodies are welded and flange welding
will start early next week.
There is continued progress in installation, the beam line in
the injector dump region is leak tight. Work continues on the
line for the recirculation dump. All remaining arc stands were
delivered and installed.
Regarding RF systems:
The 225 kV power supply was operated. We conducted a HV test
to 20 kV and ran RF from one 50 kW klystron into shorts. At close
of business Friday, work was progressing on second unit.
The elbow was affixed to the HV power supply tank for the gun.
We ran HV tests in air successfully. We are preparing for 14-15
a psi hydrostatic test early next week.
New software is ready to install for cryomodule control. We are
also preparing software changes for controlling the injector (modifications
of ITS software).
Regarding instrumentation and control systems:
We began assembling beam viewer camera mounts in tunnel. The
beam instrumentation and vacuum cabling to accelerator enclosure
was finished. Hewlett Packard workstations and X-Terminals delivered
and set-up in building. The MPS design review was held on Friday,
Sept. 12. The review committee will provide a report next week.
PSS certification for individual components continues. Buffer
boards for video multiplexing prototypes are complete and turned
over for software testing next week.
Work continued this week on the drive laser and the optical system.
After we made more modifications to the stands we are aligned
through the telescope box down to the fourth mirror can. The
two mirror cans being built by our Machine Shop were completed
today (9/12). They will be cleaned this weekend and installed
next week. The drive laser was operated about 30 hrs. this week
to continue checking out the new divide-by-40 hardware, and associated
electronics. Once again, the phase was controlled very well;
the error signal was no more than 0.4 deg. peak-to-peak. Drift
was quite low, about +/- 0.1 deg. We also continued testing
remote control of other optics equipment, and are resolving the
controls problems that have turned up.
Discussions with the two optics companies producing cavity mirrors
resulted in plans for them to send their test plates to China
Lake for testing. Rocky Mountain Instruments (RMI) sent mirrors
to Zygo to check the ROC, the results are ROC = 4.00157m, quite
close to the value determined by China Lake. Once China Lake
can tell RMI how to modify the test plate to make it correct,
they will refigure the optics. The other vendor, Research Electro-Optics
(REO) will also send China Lake their test plate, after they remove
the AR coating (which was marred by overspray) and repolish and
recoat the mirrors. They will also try annealing one mirror to
see if some of the compressive stress caused by the coating will
be relieved, which will lengthen the ROC. The right angle valve
for the upstream optical cavity assembly passed particulate and
vacuum tests, it will be installed as soon as personnel are available.
Regarding the wiggler: Checked permeability of vacuum chamber
welds. It is <1.05 which is acceptable.
Uwe Happek visited from the University of Georgia to discuss the
design and methodology of the interferometric bunch-length monitor.
He brought the first unit with him, and our plan is to acquire
and use three of them on the FEL.
Work continued toward specifying the RF phases in the injector.
Space-charge forces, because they are nonlinear and dynamic,
complicate the task of determining the transfer matrices. Our
approach is to determine the matrices for the RF elements numerically
and do multiparticle statistical simulations to study the overall
system sensitivity to RF phase. Using PARMELA is the most likely
The photocathode gun is baking in preparation for drawing electrons
off the cathode on 23 Sep 97. Numerous activities must be completed
prior to then. A partial list of activities follows, and is offered
here as an indication of the level of planning required for commissioning
a subsystem of the FEL:
Partial List of Items to Complete for Gun Turn-On
Item Person Due
Lock vacuum valve to Quarter closed. K. Jordan 9/8
Provide sketch for LCW manifold T. Siggins 9/8
Install LCW manifold for INJ N. Wilson 9/10
Install Instrument Air for HVPS purge line. J. Parkinson 9/10
Test HVPS at 15 psig with air. J. Fugitt 9/10
Install HVPS Status chassis D. Gelhaar 9/11
PSS Certifications H. Robertson 9/11
a) HVPS J. Fugitt 9/11
b) Laser Shutter K. Jordan 9/11
Fit-up HVPS tank & bellows to Gun J. Fugitt 9/15
Bake Gun & beam line. T. Siggins 9/15
Check fit-up between Gun and HVPS tank. J. Fugitt 9/15
Align Drive Laser M. Shinn 9/16
Align Light Box S. Benson 9/17
Provide Controls for: K. Jordan 9/17
a) HVPS e) Laser pulses
b) Magnet f) Telescope picomotors
c) Viewers g) Operator shutter
d) Valve h) Video & Spiricon
SOP for HVPS J. Fugitt 9/17
PSS Certification for HVPS H. Robertson 9/17
OSP for FEL building R. Legg 9/17
High Voltage process Gun T. Siggins 9/18-22
Generate 1st e-beam from Gun. R. Legg 9/23